# Using physics-inspired Singular Learning Theory to understand grokking & other phase transitions in modern neural networks

### v1

@@ 3 Experiments
The first problem setup we study are univariate polynomial regressors of increasing degree. We say a function $f:\mathbb{R}\to\mathbb{R}$ is a polynomial regressor of degree $d$ if $f(x)=\sum_{i=0}^{d}a_{i}x^{i}$ for $\{a_{i}\}_{i=0}^{d}\subset\mathbb{R}$. For a given degree $d$, we run the following procedure ten times:
@@
    Generate a $d$-degree dataset $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$ of $N=500$ samples realizable by some polynomial regressor with all coefficients in [1,1][-1,1]. We also had to enforce $x_{i}\in[-1,1]$ for numerical stability as we scaled to $d\sim 10^{3}$.
@@
to get an accurate estimate of the mean LLC $\lambda_{d}$ and its corresponding standard deviation. We present $\lambda_{d}$ for 20 choices of degree logarithmically spaced from $10^{0}$ to $10^{3}$ in Figure 5.
Figure 5: Estimated LLC versus degree of polynomial regressor (log scale), along with linear fit.
While deceptively simple, this result is actually quite interesting. Observe that polynomial regressors are regular models, meaning that there is a bijection from $\{a_{i}\}_{i=0}^{d}$ to some function from $\mathbb{R}\to\mathbb{R}$. As such, following well-established results from SLT [wei2022deep] we would expect $\lambda_{d}=\frac{d}{2}$. This means our empirical LLC estimates are lower than the theoretical expectations (note the log-scale for the degree x-axis). We reason this to be the case because the input of our polynomial regressors is not $\mathbb{R}$ but instead [1,1][-1,1], which induces parameter singularities. Explicitly stated, it might be the case that for some polynomial regression function $f_{A}$ induced by $A:=\{a_{i}\}_{i=0}^{d}\subset\mathbb{R}$ and some polynomial regression function $f_{B}$ induced by $B:=\{b_{i}\}_{i=0}^{d}\subset\mathbb{R}$, $f_{A}\mid_{[-1,1]}=f_{B}\mid_{[-1,1]}$ despite $A\neq B$. Such singularities decrease the LLC [hoogland2023weirdtrick], which is consistent with our results in Figure 5. More broadly, this shows that models on practically constrained domains can often be more singular (and thus have higher generalization capability) than expected.

### v2

@@ 3 Experiments
The first problem setup we study are univariate polynomial regressors of increasing degree. We say a function $f:\mathbb{R}\to\mathbb{R}$ is a polynomial regressor of degree $d$ if $f(x)=\sum_{i=0}^{d}a_{i}x^{i}$ for $\{a_{i}\}_{i=0}^{d}\subset\mathbb{R}$ and all $x_{i}\in\mathcal{X}\subseteq\mathbb{R}$. Note that the use of this constrained instance space $\mathcal{X}$ is required, as it is numerically unstable to use $x_{i}$ with absolute magnitude greater than one as we scale the degree $d\sim 10^{3}$. For a given degree $d$ and instance space $\mathcal{X}$, we run the following procedure ten times to get an accurate estimate of the mean LLC $\lambda_{d}$ and its corresponding standard deviation:
@@
    Generate a $d$-degree dataset $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$ of $N=500$ samples realizable by some polynomial regressor with all coefficients in [1,1][-1,1] and all samples $x_{i}\in\mathcal{X}$.
@@
We use $\mathcal{X}=[-1,1],[-0.75,0.75],$ and [0.5,0.5][-0.5,0.5] for our experiments. For each choice of $\mathcal{X}$, We present $\lambda_{d}$ for twenty choices of $d$ from $10^{0}$ to $10^{3}$ in Figure 5.
Figure 5: Estimated LLC versus degree of polynomial regressor for all choices of instance space $\mathcal{X}$.
While deceptively simple, this result is actually quite interesting. Observe that polynomial regressors are regular models, meaning that there is a bijection from $\{a_{i}\}_{i=0}^{d}$ to some function from $\mathbb{R}\to\mathbb{R}$. As such, following well-established results from SLT [wei2022deep] we would expect $\lambda_{d}=\frac{d}{2}$. This means our empirical LLC estimates, across all $\mathcal{X}$, are lower than the theoretical expectations. We reason this to be the case because the input of our polynomial regressors is not $\mathbb{R}$ but instead some choice of $\mathcal{X}\subset\mathbb{R}$, which induces parameter singularities. Restating this explicitly, it is likely that for some polynomial regression function $f_{A}$ induced by $A:=\{a_{i}\}_{i=0}^{d}\subset\mathbb{R}$ and some polynomial regression function $f_{B}$ induced by $B:=\{b_{i}\}_{i=0}^{d}\subset\mathbb{R}$, the functions $f_{A}\mid_{\mathcal{X}}=f_{B}\mid_{\mathcal{X}}$ despite $A\neq B$. Such singularities decrease the LLC [hoogland2023weirdtrick], which is consistent with our results. Furthermore, tighter interval choices for $\mathcal{X}$ would lead to more singularities and lower LLCs – this is again consistent with our results in Figure 5. This experiment establishes that models on practically constrained domains can often be more singular (and thus have higher generalization capability) than expected.